ASAP PLS

ASAP
PLS
media%2F16c%2F16c899a3-7c8e-4dc1-8d22-5c
media%2F42c%2F42c9205d-2e8a-47e5-88a6-3f
Answer #1

Solution 3:

Maturity value of bond = $1,000

Interest rate = 2.52%, 1.26% semiannual

Period to maturity = 5 years, 10 semiannual period

Bond cost today = Present value of maturity value discounted at 1.26% of 10th period

= $1,000 * PV factor at 1.26% for 10th period

= $1,000 * 0.882309

= $882.31

Solution 4:

Price per gallon of gas in past = $1.06

Price per gallon of gas today = $2.85

Time = 20 years

Let Annual rate at which per gallon cost increase every year = i

P *(1+i)^t = $2.85

$1.06 (1+i)^20 = $2.85

(1+i)^20 = $2.68867

1+i = 1.05069

i = 5.07%

Therefore annual rate at which gallon of gas increased over 20 years = 5.07%